# Probing soft signals of gravitational-wave memory with space-based interferometers

### 30 Mar 2026 v1

@@ II.3 Integrated-displacement memory
It is tempting to consider a waveform with $\frac{d^{n}h}{dt^{n}}(\infty)-\frac{d^{n}h}{dt^{n}}(-\infty)\neq 0$, with $n\geq 2$. However, such a waveform corresponds to a Riemann tensor $R_{0i0j}\approx-\frac{1}{2}\partial_{t}^{2}h_{ij}$ with a constant or polynomially growing amplitude at $t\to\pm\infty$ and is less likely to be associated with a GW [43]. Soft waveforms of the form $\frac{O}{|f|}$ and $\frac{iO}{f|f|}$ with $O\in\mathbb{R}$ are unphysical, since they correspond to time-domain waveforms that grow as $\ln|t|$ and $t\ln|t|$, respectively, as $t\to\pm\infty$.
If hereditary observables are taken into account, additional types of soft waveforms can be found. For example, the waveform $h(t)$ with a finite zeroth time moment is accompanied with an asymptotic change in the observable $O(t)\equiv\int_{-\infty}^{t}dt^{\prime}\,h(t^{\prime})$, and has the zero-frequency limit:
@@ II.4 Correction to the soft waveform
A soft waveform is the exact zero-frequency limit of the spectrum in the presence of memory. In reality, we observe the spectrum at a finite frequency, and thus will see a deviation from the soft waveform. There are no generic constraints on such corrections other than that they vanish at $f=0$, since an arbitrary waveform with zero memory can be superposed on the original waveform without affecting the memory. The correction therefore contains information from the bulk of the time interval, and the leading-order (LO) correction encodes the next-to-leading order term in the asymptotic time-domain waveform.
We focus on the case of displacement memory, and define the correction factor to be
$\mathcal{C}(f)\equiv\left(\frac{i\Delta h}{2\pi f}\right)^{-1}\tilde{h}_{\infty}(f)-1.$ (14)
For all three models in Sec. II.1, $\mathcal{C}(f)$ is analytic at $f=0$. This is generally not the case for realistic waveforms. E.g., $\lim_{t\to\pm\infty}h(t)\to-1/(2|t|)$ corresponds to $\tilde{h}_{\infty}(f)=\ln|f|$. Due to the finite signal duration and instrumental noise (the latter usually being the more limiting factor), the spectrum at $f<f_{\text{min}}$ is inaccessible. It is thus more convenient to write a generic spectrum as
$\tilde{h}_{\infty}\equiv\frac{iHe^{-2i\varsigma}}{2\pi f}[1+C(f)],$ (15)
with $H>0$, $\varsigma\in\mathbb{R}$, and $C(f_{\text{min}})=0$. As $f_{\text{min}}\to 0$ (which requires the measurement time $T\to\infty$), $\{\varsigma,H\}\to\{0,\Delta h\}$, if the memory truly persists. Other reasonable definitions of $H$ are possible, but we are interested in situations where $\Delta h$ is not too different from $i2\pi f_{\text{min}}\tilde{h}_{\infty}(f_{\text{min}})$ (otherwise the soft waveform would be a poor approximation) so the corresponding correction factors $C(f)\equiv\left(\frac{iH}{2\pi f}\right)^{-1}\tilde{h}_{\infty}(f)-1$ should not differ significantly. Note that $C(f)$ cannot generally be obtained from the Taylor expansion about $f_{\text{min}}$, which may not converge away from $f_{\text{min}}$. If both linear polarization components are described by soft waveforms of the same type, and their arrival times $t_{*}^{\lambda}$ satisfies $f|(t_{*}^{+}-t_{*}^{\times})|\ll 1$, then the GW is approximately linearly polarized (i.e., $\tilde{h}_{+}/\tilde{h}_{\times}\in\mathbb{R}$). This may not hold in general due to finite-frequency corrections. Nevertheless, we define $\tilde{h}_{+}-i\tilde{h}_{\times}=\frac{iH_{+}e^{-2i\varsigma_{+}}}{2\pi f}[1+C_{+}(f)]+\frac{H_{\times}e^{-2i\varsigma_{\times}}}{2\pi f}[1+C_{\times}(f)]\equiv\frac{iHe^{-2i\psi}}{2\pi f}[1+C(f)]$, where $H$ and $C(f)$ are the effective memory amplitude and correction factor, respectively.
@@ III Memory of compact binaries
$h(u=t-R)=h_{+}-ih_{\times}\equiv\sum_{l\geq 2,m}h_{l,m}(u)\,_{-2}Y_{lm}(\Theta,\Phi),$ (16)
with $R$ being the observer’s distance to the source, $(\Theta,\Phi)$ the spherical polar coordinates of the observer, and $_{s}Y_{lm}$ the spin-weighted spherical harmonics with spin weight $s$. Here the TT polarization is defined by $\mathbf{a}=\mathbf{e}_{\Theta}$ and $\mathbf{b}=\mathbf{e}_{\Phi}$ in Eq. (3). We also use the Cartesian coordinates $(X,Y,Z)=R(s_{\Theta}c_{\Phi},s_{\Theta}s_{\Phi},c_{\Theta})$, where $c_{z}\equiv\cos z$ and $s_{z}\equiv\sin z$.
@@
III.1 Propagation of waveforms
Eq. (16) is an idealized description, since in reality, the spacetime is not asymptotically flat. For a locally plane GW with sufficiently high frequency, [^1] Here we consider a frequency band that is low enough for the soft waveform to be a good approximation, while still high enough for the eikonal approximation to be applicable. its propagation in a curved spacetime background can be described by the eikonal approximation. If the observed frequency of a Fourier mode with original frequency $f$ is shifted to $f^{\prime}=\varrho f$, the observed frequency spectrum $\tilde{h}_{\infty}^{\prime}(f)$ is related to the original spectrum $\tilde{h}_{\infty}(f)$ via
@@ III.1 Propagation of waveforms
The zeroth order eikonal approximation of massless relativistic waves due to a curved spacetime background is captured by the massless Klein-Gordon equation in the background coordinates. We consider a spatially flat Friedmann-Lemaitre-Robertson-Walker spacetime that is weakly perturbed: ${ds^{2}}/{a^{2}}=-(1+2\Psi)d\eta^{2}+2\Xi_{i}d\eta dx^{i}+(\delta_{ij}+H_{ij})dx^{i}dx^{j}$, where $a$ is the cosmological scale factor, $d\eta\equiv dt/a$ is the conformal time. The massless Klein-Gordon equation in the coordinates $(\eta,\mathbf{x})$ reads ${\partial_{a}\left(\sqrt{-g}\,g^{ab}\partial_{b}\phi\right)}/{\sqrt{-g}}=(W+B^{a}\partial_{a})\phi=0$, where $W\equiv g^{ab}\partial_{a}\partial_{b}$ and $B^{a}\equiv\partial_{b}(\sqrt{-g}\,g^{ab})/\sqrt{-g}$. In the eikonal approximation [90], the complex scalar field is described by a rapidly varying real phase $S(\eta,\mathbf{x})$ and a slowly varying real amplitude $A(\eta,\mathbf{x})$ as $\phi=Ae^{iS}$, with $\omega(\eta,\mathbf{x})\equiv-\partial_{\eta}S=2\pi f$ and $\mathbf{k}(\eta,\mathbf{x})\equiv\nabla S$. At zeroth order, $W\phi\approx 0$, and $k_{a}=\partial_{a}S=(-\omega,\mathbf{k})$ satisfies the local dispersion relation $D(x,k)\approx-g^{ab}k_{a}k_{b}\approx(1-2\Psi)\omega^{2}-|\mathbf{k}|^{2}+H_{ij}k_{i}k_{j}+2\Xi_{i}\omega k_{i}=0$. $k_{a}(\eta,\mathbf{x}(\eta))$ evolves according to ${dk_{a}}/{d\eta}\approx(\partial_{a}D)/(\partial_{\omega}D)$, e.g.,
@@
Meanwhile, the amplitude evolves according to $\left[(\partial_{k_{a}}D)\partial_{a}+\left(\partial_{k_{a}}\partial_{k_{b}}D\right)(\partial_{a}k_{b})/2-B^{a}k_{a}\right]A\approx 0$, i.e., $k^{a}\partial_{a}\ln A\approx-\nabla_{a}k^{a}/2$. The frequency measured by a timelike observer with four-velocity $u^{\mu}$ (normalized as $u^{\mu}u_{\mu}=-1$) is then $\omega_{\text{o}}=-u^{\mu}k_{\mu}$. E.g., for an observer (emitter) with $u_{\text{o(e)}}^{\mu}\approx a_{\text{o(e)}}^{-1}[1-\Psi_{\text{o(e)}}]\delta^{\mu}_{0}$ at the time $\eta_{\text{o(e)}}$ of observation (emission), $\omega_{\text{o(e)}}\approx a_{\text{o(e)}}^{-1}[1-\Psi_{\text{o(e)}}]\,\omega(\eta_{\text{o(e)}})$. In the absence of metric perturbations, $\omega(\eta_{\text{o}})=\omega(\eta_{\text{e}})$, thus $\omega_{\text{o}}=\omega_{\text{e}}/(1+z)$ with $z\equiv a_{\text{o}}/a_{\text{e}}-1$ being the redshift of the source.
@@ III.2 Displacement memory from compact binary scattering
The displacement and velocity memory can appear at 0PN in a linear gravitational radiation, since according to the quadrupole formula (17), they can be produced respectively by the asymptotic changes of the second and third time derivatives of $M_{ij}$:
$\ddot{M}_{ij}$ $=\int d^{3}x\,\rho\,(2v_{i}v_{j}+a_{i}x_{j}+x_{i}a_{j}),$ (21)
$\dddot{M}_{ij}$ $=\int d^{3}x\,\rho\,(3a_{i}v_{j}+3v_{i}a_{j}+\dot{a}_{i}x_{j}+x_{i}\dot{a}_{j}),$
@@
where $v_{i}={dx^{i}}/{dt}$ is the velocity and $a_{i}={dv^{i}}/{dt}$ is the acceleration of the nonrelativistic matter. Velocity memory can thus be produced by an asymptotic change of acceleration accompanied by nearly constant velocity over the relevant time scale [43, 9], whereas displacement memory is produced in generic scattering processes. [^2] Similarly, the scattering of scalar or vector charges can also produce memory in the corresponding radiation (see Appendix B for a brief discussion). In this section, we examine the soft waveforms of compact binary scattering events in the moderately relativistic regime, where the PN approximation is applicable. We describe the scattering in the two-body center-of-mass (CM) frame, and denote $x_{I}^{\mu}=(t,\mathbf{x}_{I}(t))$ the 4-position of the $I$-th body with mass $m_{I}$. We also define $\mathbf{x}\equiv\mathbf{x}_{2}-\mathbf{x}_{1}$, $\mathbf{v}\equiv\dot{\mathbf{x}}$ and $r=|\mathbf{x}|$. In the CM frame, the initial state is fully specified by the relative velocity $\mathbf{v}_{0}=\mathbf{v}(t\to-\infty)$ and the impact parameter vector $\mathbf{b}=[\mathbf{r}-(\mathbf{r}\cdot\hat{\mathbf{v}})\,\hat{\mathbf{v}}]_{t\to-\infty}$, with $\hat{\mathbf{v}}=\mathbf{v}/|\mathbf{v}|$; we denote $v_{0}=|\mathbf{v}_{0}|$ and $b=|\mathbf{b}|$. Covariantly, the initial state can be defined by the four-velocities $u_{I}^{\mu}=(1-|\mathbf{v}_{I}|^{2})^{-1/2}(1,\mathbf{v}_{I})$ with an impact parameter four-vector $b^{\mu}$ satisfying $b_{\mu}u_{1}^{\mu}=b_{\mu}u_{2}^{\mu}=0$. The CM frame is defined by $m_{1}u_{1}^{i}+m_{2}u_{2}^{i}=0$, in which $b^{\mu}=(0,\mathbf{b})$.
@@
We consider first the quadrupole waveform of a Newtonian hyperbolic encounter. The orbit of $\mathbf{x}(t)$ is taken to be a Keplerian orbit with semi-major axis $a$ and eccentricity $e$. Correspondingly, $b=a\sqrt{e^{2}-1}$, $v_{0}=\sqrt{M/a}$; so $e=\sqrt{1+(b/M)^{2}v_{0}^{4}}$, $a=M/v_{0}^{2}$. This approximation is reasonable if the incident velocity $v_{0}$ (for a given impact parameter $b$) is neither too large nor too small, since in the former case the conservative PN corrections become important, while in the latter case the gravitational radiation reaction matters. [^3] For example, a Keplerian parabolic orbit produces only null displacement memory [32]. However, when gravitational radiation reaction is included, a final state with vanishing relative velocity arises from nonzero initial relative velocity, leading to nonzero ordinary displacement memory. The quadrupole moment in the CM frame is given by $M_{ij}=\nu Mx_{i}x_{j}$, where $M=m_{1}+m_{2}$ is the total mass and $\nu=m_{1}m_{2}/M^{2}\in(0,1/4]$ is the symmetric mass ratio.
@@
We choose a source frame in which the orbital angular momentum aligns with the $Z$-axis, with the periastron on the $X$ axis; we also shift the time coordinate such that $t=0$ corresponds to the periastron crossing. With this choice, only $h_{2,\pm 2}$ and $h_{2,0}$ are non-vanishing. Using the eccentric-anomaly parameterization of the hyperbolic orbit (see, e.g., [15]) and the integral representation of the Hankel function of the first kind: $H_{m}^{(1)}(z)=\frac{1}{i\pi}\int_{-\infty}^{\infty}d\xi\,e^{z\sinh\xi-m\xi}$, the idealized spectrum of the quadrupole waveform can be derived as
$\frac{f\tilde{h}_{2,\pm 2}/\varpi}{2\pi h_{0}}$ $=\left(\frac{\pm\sqrt{e^{2}-1}}{2e}+\frac{e^{2}-1}{e}\varpi\right)H_{1+in}$ (22)
@@
$\frac{f\tilde{h}_{2,0}}{h_{0}}=-\frac{i\pi}{\sqrt{6}}\varpi H_{in},$ (23)
with $h_{0}\equiv\sqrt{{64}/{(5\pi)}}\,{\nu M^{2}}/{(aR)}$, $\varpi\equiv\pi f/\Omega$, $\Omega\equiv\sqrt{M/a^{3}}$, $n\equiv 2\varpi$, and $H_{m}\equiv H_{m}^{(1)}(ine)$. The relevant spin-weighted spherical harmonics are
$\left(\begin{smallmatrix}{}_{-2}Y_{2,2}\\ {}_{-2}Y_{2,0}\\ {}_{-2}Y_{2,-2}\end{smallmatrix}\right)=\sqrt{\frac{5}{4\pi}}\left(\begin{smallmatrix}e^{2i\Phi}c^{4}_{\Theta/2}\\ \sqrt{\frac{3}{8}}s_{\Theta}^{2}\\ e^{-2i\Phi}s^{4}_{\Theta/2}\end{smallmatrix}\right).$ (24)
Using the limits: $\lim_{n\to 0}H_{in}\to i\left(\frac{2}{\pi}\ln ne\right)$ and $\lim_{n\to 0}H_{1+in}\to\left(-\frac{2}{\pi en}-\frac{1}{e}\right)+i\left(\frac{2}{\pi e}\ln ne\right)$, we obtain from Eqs. (22)-(23) the soft waveform $\tilde{h}(f)\to{i(\Delta h_{+}-i\Delta h_{\times})}/{(2\pi f)}$ (since the periastron crossing happens at $t=t_{*}=0$) with its LO correction:
$\frac{f\tilde{h}_{2,\pm 2}}{h_{0}}\approx$ $\mp\frac{\sqrt{e^{2}-1}}{e^{2}}(1+\pi\varpi)+2\left(\frac{1}{e^{2}}-1\right)\varpi$ (25)

### 13 Apr 2026 v2

@@ II.3 Integrated-displacement memory
It is tempting to consider a waveform with $\frac{d^{n}h}{dt^{n}}(\infty)-\frac{d^{n}h}{dt^{n}}(-\infty)\neq 0$, with $n\geq 2$. However, such a waveform corresponds to a Riemann tensor $R_{0i0j}\approx-\frac{1}{2}\partial_{t}^{2}h_{ij}$ with a constant or polynomially growing amplitude at $t\to\pm\infty$ and is less likely to be associated with a GW [48]. Waveforms of the form $\frac{O}{|f|}$ and $\frac{iO}{f|f|}$ with $O\in\mathbb{R}$ are unphysical, since they correspond to time-domain waveforms that grow as $\ln|t|$ and $t\ln|t|$, respectively, as $t\to\pm\infty$.
If hereditary observables are taken into account, additional types of soft waveforms can be defined. For example, the waveform $h(t)$ with a finite zeroth time moment is accompanied with an asymptotic change in the observable $O(t)\equiv\int_{-\infty}^{t}dt^{\prime}\,h(t^{\prime})$, and has the zero-frequency limit:
@@ II.4 Correction to the soft waveform
A soft waveform is the exact zero-frequency limit of the spectrum in the presence of memory. In reality, we observe the spectrum at a finite frequency, and thus will see a deviation from the soft waveform. There are no generic constraints on such corrections other than that they vanish at $f=0$, since an arbitrary waveform $\mathtt{h}(t)$ with zero memory can be superposed on the original waveform without affecting the memory. The correction therefore contains information from the bulk of the time interval, and the leading-order (LO) correction encodes the next-to-leading order term in the asymptotic time-domain waveform.
We focus on the case of displacement memory (setting $t_{*}=0$), and define the exact correction factor to be
$\mathcal{C}(f)\equiv\left(\frac{2\pi f}{i\Delta h}\right)\tilde{h}_{\infty}(f)-1=[\mathcal{C}(-f)]^{*}.$ (12)
For all three models in Sec. II.1, $\mathcal{C}(f)$ is analytic [^2] For the arctan model, $\lim_{f\to 0}\mathcal{C}(f)=(-2\pi/f_{*})|f|$; for the tanh model, $\lim_{f\to 0}\mathcal{C}(f)=[-\pi^{4}/(6f_{*}^{2})]f^{2}$; for the linear model, $\lim_{f\to 0}\mathcal{C}(f)=i(\pi/f_{*})f$. at $f=0$. This is generally not the case for realistic waveforms. E.g., $\lim_{|t|\to\infty}\mathtt{h}(t)\to-1/(2|t|)$ corresponds to $\tilde{\mathtt{h}}_{\infty}(f\to 0)=\ln|f|$; $\lim_{|t|\to\infty}\mathtt{h}(t)\propto|t|^{-\alpha}$ with $\alpha\in(0,1)$ corresponds to $\tilde{\mathtt{h}}_{\infty}(f\to 0)\propto|f|^{\alpha-1}$. Due to the finite signal duration and instrumental noise (the latter usually being the more limiting factor), the spectrum at sufficiently low frequencies is inaccessible. It can thus be useful to introduce a finite frequency $f_{\text{min}}$ and write
$\tilde{h}_{\infty}\equiv\frac{iH}{2\pi f}[1+C(f)],$ (13)
with $H\in\mathbb{C}$ and $C(f_{\text{min}})=0$. As $f_{\text{min}}\to 0$ (which requires the measurement time $T\to\infty$), $H\to\Delta h$ and $C(f)\to\mathcal{C}(f)$, if the memory truly persists. [^3] In reality, since the distance between the observer and the GW source is finite, the displacement memory should vanish as $t\to\infty$ [99]. However, $\mathcal{C}(f)$ predicted by the waveform in the distant-source approximation is expected to be observable before reaching the frequency corresponding to such a long time scale. Note that $C(f)$ cannot generally be obtained from the Taylor expansion about $f_{\text{min}}$, which may not converge away from $f_{\text{min}}$. We are interested in situations where both linear polarization components can be approximately described by soft waveforms of the same type, and their arrival times $t_{*}^{\lambda}\approx t_{*}$ satisfies $f|(t_{*}^{+}-t_{*}^{\times})|\ll 1$. In the zero-frequency limit, the GW is then linearly polarized (i.e., $\tilde{h}_{+}/\tilde{h}_{\times}\in\mathbb{R}$). This may not hold in general due to finite-frequency corrections. Nevertheless, we can define $\tilde{h}_{+}-i\tilde{h}_{\times}\equiv\frac{i}{2\pi f}(\Delta h_{+}-i\Delta h_{\times})e^{2\pi if_{*}t}[1+\mathcal{C}(f)]\equiv\frac{iHe^{-2i\psi}}{2\pi f}e^{2\pi if_{*}t}[1+C(f)]$, with $\psi\in\mathbb{R}$, where $H>0$ is the effective memory amplitude and $C(f)$ is the correction factor, defined on the two-sided frequency domain.
@@ III Memory of compact binaries
$h(u)=h_{+}-ih_{\times}\equiv\sum_{l\geq 2,m}h_{l,m}(u)\,_{-2}Y_{lm}(\Theta,\Phi),$ (14)
where u=tRu=t-R, $R$ is the observer’s distance to the source, $(\Theta,\Phi)$ are the spherical polar coordinates of the observer, and $_{s}Y_{lm}$ are the spin-weighted spherical harmonics with spin weight $s$. Here the TT polarization is defined by $\mathbf{a}=\mathbf{e}_{\Theta}$ and $\mathbf{b}=\mathbf{e}_{\Phi}$ in Eq. (3). We also use the Cartesian coordinates $(X,Y,Z)=R(s_{\Theta}c_{\Phi},s_{\Theta}s_{\Phi},c_{\Theta})$, where $c_{z}\equiv\cos z$ and $s_{z}\equiv\sin z$.
@@
III.1 Waveform propagation
Eq. (14) is an idealized description, since in reality the spacetime is not asymptotically flat and the observer is not located at infinity. For a locally plane GW with sufficiently high frequency, its propagation in a curved spacetime background can be described by the eikonal approximation. If the observed frequency of a Fourier mode with original frequency $f$ is shifted to $f^{\prime}=\varrho f$, the observed frequency spectrum $\tilde{h}_{\infty}^{\prime}(f)$ is related to the original spectrum $\tilde{h}_{\infty}(f)$ via
@@ III.1 Waveform propagation
The zeroth order eikonal approximation of massless relativistic waves in a curved spacetime background is captured by the massless Klein-Gordon equation in the background coordinates. We consider a spatially flat Friedmann-Lemaitre-Robertson-Walker spacetime that is weakly perturbed: ${ds^{2}}/{a^{2}}=-(1+2\Psi)d\eta^{2}+2\Xi_{i}d\eta dx^{i}+(\delta_{ij}+H_{ij})dx^{i}dx^{j}$, where $a(\eta)$ is the cosmological scale factor, $d\eta\equiv dt/a$ is the conformal time. The massless Klein-Gordon equation in the coordinates $(\eta,\mathbf{x})$ reads ${\partial_{a}\left(\sqrt{-g}\,g^{ab}\partial_{b}\phi\right)}/{\sqrt{-g}}=(W+B^{a}\partial_{a})\phi=0$, where $W\equiv g^{ab}\partial_{a}\partial_{b}$ and $B^{a}\equiv\partial_{b}(\sqrt{-g}\,g^{ab})/\sqrt{-g}$. In the eikonal approximation [97], a Fourier mode of the complex scalar field is described by a rapidly varying real phase $S(\eta,\mathbf{x})$ and a slowly varying real amplitude $A(\eta,\mathbf{x})$ as $\phi=Ae^{iS}$, with $\omega(\eta,\mathbf{x})\equiv-\partial_{\eta}S=2\pi f$ and $\mathbf{k}(\eta,\mathbf{x})\equiv\partial_{\mathbf{x}}S$. At zeroth order, $W\phi\approx 0$, thus $k_{\mu}=\partial_{\mu}S=(-\omega,\mathbf{k})$ satisfies the local dispersion relation: $0=D(x,k)=-g^{ab}k_{a}k_{b}\approx a^{-2}[(1-2\Psi)\omega^{2}-|\mathbf{k}|^{2}+H_{ij}k_{i}k_{j}+2\Xi_{i}\omega k_{i}]$. $k_{\mu}(\eta,\mathbf{x}(\eta))$ evolves according to ${dk_{\mu}}/{d\eta}=(\partial_{\mu}D)/(\partial_{\omega}D)$, e.g.,
@@
Meanwhile, the amplitude evolves according to $\left[(\partial_{k_{a}}D)\partial_{a}+\left(\partial_{k_{a}}\partial_{k_{b}}D\right)(\partial_{a}k_{b})/2-B^{a}k_{a}\right]A\approx 0$, i.e., $k^{a}\partial_{a}\ln A\approx-\nabla_{a}k^{a}/2$. The frequency measured by a timelike observer with four-velocity $u^{\mu}$ (normalized as $u^{\mu}u_{\mu}=-1$) is $\omega_{\text{o}}=-u^{\mu}k_{\mu}$. E.g., for an observer (emitter) with $u_{\text{o(e)}}^{\mu}\approx a_{\text{o(e)}}^{-1}[1-\Psi_{\text{o(e)}}]\delta^{\mu}_{0}$ at the time $\eta_{\text{o(e)}}$ of observation (emission), $\omega_{\text{o(e)}}\approx a_{\text{o(e)}}^{-1}[1-\Psi_{\text{o(e)}}]\,\omega(\eta_{\text{o(e)}})$. In the absence of metric perturbations, $\omega(\eta_{\text{o}})=\omega(\eta_{\text{e}})$, thus $\omega_{\text{o}}=\omega_{\text{e}}/(1+z)$ with $z\equiv a_{\text{o}}/a_{\text{e}}-1$ being the redshift of the source.
@@ III.2 Displacement memory from compact binary scattering
The displacement and velocity memory can appear at 0PN in linearized gravitational radiation, since according to the quadrupole formula (15), they can be produced respectively by the asymptotic changes of the second and third time derivatives of $M_{ij}$, namely $\ddot{M}_{ij}=\int d^{3}x\,\rho\,(2v_{i}v_{j}+a_{i}x_{j}+x_{i}a_{j})$ and $\dddot{M}_{ij}=\int d^{3}x\,\rho\,(3a_{i}v_{j}+3v_{i}a_{j}+\dot{a}_{i}x_{j}+x_{i}\dot{a}_{j})$, where $v_{i}={dx^{i}}/{dt}$ is the velocity and $a_{i}={dv^{i}}/{dt}$ is the acceleration of the nonrelativistic matter. Velocity memory can thus be produced by an asymptotic change of acceleration accompanied by nearly constant velocity over the relevant time scale [48, 11], whereas displacement memory is produced in generic scattering processes. [^4] In an asymptotically flat spacetime, the displacement memory at future null infinity can be decomposed into ordinary and null components [10], sourced by the energy–momentum of massive matter and null energy flux, respectively. In this section, we examine the soft waveforms of compact binary scattering events in the moderately relativistic regime, where the PN approximation is applicable. We describe the scattering in the two-body center-of-mass (CM) frame, and denote $x_{I}^{\mu}=(t,\mathbf{x}_{I}(t))$ the 4-position of the $I$-th body with mass $m_{I}$. We also define $\mathbf{x}\equiv\mathbf{x}_{2}-\mathbf{x}_{1}$, $\mathbf{v}\equiv\dot{\mathbf{x}}$ and $r=|\mathbf{x}|$. In the CM frame, the initial state can be specified by the relative velocity $\mathbf{v}_{0}=\mathbf{v}(t\to-\infty)$ and the impact parameter vector $\mathbf{b}=[\mathbf{r}-(\mathbf{r}\cdot\hat{\mathbf{v}})\,\hat{\mathbf{v}}]_{t\to-\infty}$, with $\hat{\mathbf{v}}=\mathbf{v}/|\mathbf{v}|$; we denote $v_{0}=|\mathbf{v}_{0}|$ and $b=|\mathbf{b}|$. Covariantly, the initial state can be defined by the four-velocities $u_{I}^{\mu}=(1-|\mathbf{v}_{I}|^{2})^{-1/2}(1,\mathbf{v}_{I})$ with an impact parameter four-vector $b^{\mu}$ satisfying $b_{\mu}u_{1}^{\mu}=b_{\mu}u_{2}^{\mu}=0$. The CM frame is defined by $m_{1}u_{1}^{i}+m_{2}u_{2}^{i}=0$, in which $b^{\mu}=(0,\mathbf{b})$.
@@
We first consider the 0PN quadrupole waveform of a Newtonian hyperbolic encounter. The orbit of $\mathbf{x}(t)$ is taken to be a Keplerian orbit with semi-major axis $a$ and eccentricity $e$. Correspondingly, $b=a\sqrt{e^{2}-1}$, $v_{0}=\sqrt{M/a}$; so $e=\sqrt{1+(b/M)^{2}v_{0}^{4}}$, $a=M/v_{0}^{2}$. This approximation is reasonable if the incident velocity $v_{0}$ (for a given impact parameter $b$) is neither too large nor too small, since in the former case the conservative PN corrections become important, while in the latter case the gravitational radiation reaction matters. [^5] For example, a Keplerian parabolic orbit produces only null displacement memory [35]. However, when gravitational radiation reaction is included, a final state with vanishing relative velocity arises from nonzero initial relative velocity, leading to nonzero ordinary displacement memory. The quadrupole moment in the CM frame is given by $M_{ij}=\nu Mx_{i}x_{j}$, where $M=m_{1}+m_{2}$ is the total mass and $\nu=m_{1}m_{2}/M^{2}\in(0,1/4]$ is the symmetric mass ratio.
@@
We choose a source frame in which the orbital angular momentum is aligned with the $Z$-axis and the periastron lies on the $X$-axis; the time coordinate is shifted such that $t=0$ corresponds to the periastron crossing. With this choice, only $h_{2,\pm 2}$ and $h_{2,0}$ are nonzero. Using the eccentric-anomaly parameterization of the hyperbolic orbit (see, e.g., [18]) and the integral representation of the Hankel function of the first kind: $H_{m}^{(1)}(z)=\int_{-\infty}^{\infty}\frac{d\xi}{i\pi}e^{z\sinh\xi-m\xi}$, the idealized positive-frequency spectrum of the 0PN waveform can be derived as
$\frac{f\tilde{h}_{2,\pm 2}(f)/\varpi}{2\pi h_{0}}$ $=\left(\frac{\pm\sqrt{e^{2}-1}}{2e}+\frac{e^{2}-1}{e}\varpi\right)H_{1+in}$ (19)
@@
$\frac{f\tilde{h}_{2,0}(f)}{h_{0}}=-\frac{i\pi}{\sqrt{6}}\varpi H_{in},$ (20)
with $h_{0}\equiv\sqrt{{64}/{(5\pi)}}\,{\nu M^{2}}/{(aR)}$, $\varpi\equiv\pi f/\Omega$, $\Omega\equiv\sqrt{M/a^{3}}$, $n\equiv 2\varpi$, and $H_{m}\equiv H_{m}^{(1)}(ine)$. Note that $\tilde{h}_{2,2}(-f)=[\tilde{h}_{2,-2}(f)]^{*}$, $\tilde{h}_{2,0}(-f)=[\tilde{h}_{2,0}(f)]^{*}$. Using the limits: $\lim_{n\to 0}H_{in}\to i\left(\frac{2}{\pi}\ln ne\right)$ and $\lim_{n\to 0}H_{1+in}\to\left(-\frac{2}{\pi en}-\frac{1}{e}\right)+i\left(\frac{2}{\pi e}\ln ne\right)$, we obtain from Eqs. (19)-(20) the soft waveform $\tilde{h}(f)\to{i(\Delta h_{+}-i\Delta h_{\times})}/{(2\pi f)}$ (since the periastron crossing happens at $t=t_{*}=0$) with its LO correction:
$\frac{f\tilde{h}_{2,\pm 2}(f)}{h_{0}}\approx$ $\mp\frac{\sqrt{e^{2}-1}}{e^{2}}(1+\pi\varpi)+2\left(\frac{1}{e^{2}}-1\right)\varpi$ (21)